OpenStudy (anonymous):

Multiply: x^2-3x-10/x^2-6x+5 times x-1/x^2-4

OpenStudy (anonymous):

Factored: \[\frac{ (x-5)(x+7) }{(x-5)(x-1)} x \frac{ x-1 }{ (x+2)(x-2) }\]

OpenStudy (anonymous):

Oops first part ^^ (x-5)(x+2)

OpenStudy (anonymous):

Crossed out: 1/(x-2) Is my answer right?

jimthompson5910 (jim_thompson5910):

hopefully you got this \[\large \frac{ (x-5)(x+2) }{(x-5)(x-1)} x \frac{ x-1 }{ (x+2)(x-2) }\] after you factored everything

OpenStudy (anonymous):

Yup.

jimthompson5910 (jim_thompson5910):

you are correct, it's \[\Large \frac{1}{x-2}\]

OpenStudy (anonymous):

Okay thanks for confirming . :)

jimthompson5910 (jim_thompson5910):

yw